OpenStudy (anonymous):

Subtract ((x-3)/(3))-((x-4)/(4))

OpenStudy (binary3i):

\[((x-3)/3)-((x-4)/4)=(x/3)-(3/3)-((x/4)-(4/4))=(x/3)-1-(x/4)+1=(x/3)-(x/4)\]

OpenStudy (binary3i):

so it becomes \[(x/3)-(x/4)\]

OpenStudy (anonymous):

common denominator, so 4 is denominator on the right and 3 on the left, multiply left by 4 and right by 3, making common denominator 12....... so then you get 4*(x-3) over 12 and 3*(x-4) over 12 4x-12-3x-(-12) over 12 or x/12

OpenStudy (anonymous):

Thank you!